# Charged superradiant instability in a spherical regular black hole

### v3

@@ Abstract
We examine the stability of a spherically symmetric regular black hole when subjected to perturbations from a charged scalar field. This particular black hole is constructed by deforming the Minkowski spacetime. It has been observed that the charged superradiant instability arises only within a specific range of the deformation parameter, potentially resulting in an instability growth rate with a maximum magnitude of approximately $\text{Im}(M\omega)\sim 10^{-3}$. This growth rate significantly exceeds the instability identified in ABG black holes discussed in prior research, suggesting a notable timescale for detecting this phenomenon in astrophysical scenarios. Additionally, we conduct a thorough investigation into how the three parameters of the model influence the onset and intensity of the instability. Our analysis offers further insights into the possible emergence of this instability in spherically regular black holes and its association with the nonlinear effects of the electromagnetic field.
@@ I Introduction
Singularities are commonly considered non-physical within the classical theoretical framework due to their infinite energy density, prompting researchers to seek resolution through the inclusion of quantum effects in matter or gravity. While singularity theorems rely on strict conditions like the null energy condition for matter, allowing for the possibility of singularity-free black holes. To address the issue of non-physical singularities, a significant effort is being made by researchers to develop regular black hole solutions. This endeavor dates back to 1968 when Bardeen introduced a regular black hole model that adhered to the weak energy condition [14], circumventing the singularity theorems. Subsequently, Ayon-Beato and Garcia established a class of regular black holes known as ABG black holes [15, 16], featuring spherical charged configurations within the framework of Einstein-nonlinear electromagnetic fields. Further advancements led them to realize the Bardeen black hole as an exact solution within some Einstein-nonlinear-Maxwell theory [17]. These breakthroughs have inspired the development of various regular black hole models predominantly supported by nonlinear electromagnetic fields (see for example [18, 19, 20, 21, 22, 23] and references therein), which serve as an avenue to incorporate quantum effects in the electromagnetic field theory [24, 25, 26, 27]. Extensive investigations have been conducted on the dynamic behaviors, thermodynamic properties, and astrophysical implications of these black holes. The distinct observational effects exhibited by regular black holes compared to conventional ones have been a subject of interest, as indicated by recent studies and reviews on the topic [20, 21, 22, 28, 29, 30, 31].
@@
The authors of Refs. [32, 33, 34] recently conducted an extensive examination of the superradiance phenomenon in ABG black holes subjected to charged scalar field perturbations. Their findings suggest that within certain parameter ranges, ABG black holes display superradiant instability, a characteristic distinct from the well-established behavior of Reissner-Nordström (RN) black holes. The study of superradiance and its associated instability in black holes has been a focal point of research in black hole physics for a significant period. Superradiant instability typically manifests in rotating black holes [35, 36, 37, 38, 39]. However, for non-rotating, spherically symmetric RN black holes, superradiance can occur under charged scalar field perturbations but leading no instability [40, 41], unless the spacetime has a special boundary or an artificial boundary condition is imposed [42, 43, 44, 45, 46, 47]. The research outcomes in the Refs. [32, 33, 34] introduce a novel scenario: when accounting for the nonlinear effects of the electromagnetic field, spherically symmetric charged regular black holes may also exhibit superradiant instability. The superradiant instability of black holes holds significant astronomical observational implications. This mechanism results in the accumulation of scalar particles around the black hole, giving rise to observable effects such as alterations in gravitational waveforms and interference with the imaging of the black hole shadow [48, 49, 50, 51, 52, 53, 54]. These observable effects offer valuable insights and avenues for exploring exotic particles beyond the standard model and furthering our comprehension of the physical characteristics of black holes. See more discussions on this phenomenon and its astrophysical implications, please refer to the review [55] and references therein. Thus, the research findings in Refs. [32, 33, 34] not only unveil the superradiant instability of ABG black holes under specific circumstances but also present a potential observational approach for identifying and investigating the properties of regular black holes.

### v4

@@ Abstract
We examine the stability of a spherically symmetric regular black hole when subjected to perturbations from a charged scalar field. This particular black hole is constructed by deforming the Minkowski spacetime. It has been observed that the charged superradiant instability arises only within a specific range of the deformation parameter, potentially resulting in an instability growth rate with a maximum magnitude of approximately $\text{Im}(M\omega)\sim 10^{-3}$. This growth rate significantly exceeds the instability identified in Ayón-Beato-García (ABG) black holes discussed in prior research, suggesting a notable timescale for detecting this phenomenon in astrophysical scenarios. Additionally, we conduct a thorough investigation into how the three parameters of the model influence the onset and intensity of the instability. Our analysis offers further insights into the possible emergence of this instability in spherical regular black holes and its association with the nonlinear effects of the electromagnetic field.
@@ I Introduction
Singularities are commonly considered non-physical within the classical theoretical framework due to their infinite energy density, prompting researchers to seek resolution through the inclusion of quantum effects in matter or gravity. While singularity theorems rely on strict conditions like the null energy condition for matter, allowing for the possibility of singularity-free black holes. To address the issue of non-physical singularities, a significant effort is being made by researchers to develop regular black hole solutions. This endeavor dates back to 1968 when Bardeen introduced a regular black hole model that adhered to the weak energy condition [14], circumventing the singularity theorems. Subsequently, Ayón-Beato and García established a class of regular black holes known as ABG black holes [15, 16], featuring spherical charged configurations within the framework of Einstein-nonlinear electromagnetic fields. Further advancements led them to realize the Bardeen black hole as an exact solution within some Einstein-nonlinear-Maxwell theory [17]. These breakthroughs have inspired the development of various regular black hole models predominantly supported by nonlinear electromagnetic fields (see for example [18, 19, 20, 21, 22, 23] and references therein), which serve as an avenue to incorporate quantum effects in the electromagnetic field theory [24, 25, 26, 27]. Extensive investigations have been conducted on the dynamic behaviors, thermodynamic properties, and astrophysical implications of these black holes. The distinct observational effects exhibited by regular black holes compared to conventional ones have been a subject of interest, as indicated by recent studies and reviews on the topic [28, 29, 20, 21, 22, 30, 31, 32, 33, 34].
@@
The authors of Refs. [35, 36, 37] recently conducted an extensive examination of the superradiance phenomenon in ABG black holes subjected to charged scalar field perturbations. Their findings suggest that within certain parameter ranges, ABG black holes display superradiant instability, a characteristic distinct from the well-established behavior of Reissner-Nordström (RN) black holes. The study of superradiance and its associated instability in black holes has been a focal point of research in black hole physics for a significant period. Superradiant instability typically manifests in rotating black holes [38, 39, 40, 41, 42, 43]. However, for non-rotating, spherically symmetric RN black holes, superradiance can occur under charged scalar field perturbations but leading no instability [44, 45], unless the spacetime has a special boundary or an artificial boundary condition is imposed [46, 47, 48, 49, 50, 51, 52, 53]. The research outcomes in the Refs. [35, 36, 37] introduce a novel scenario: when accounting for the nonlinear effects of the electromagnetic field, spherically symmetric charged regular black holes may also exhibit superradiant instability. The superradiant instability of black holes holds significant astronomical observational implications. This mechanism results in the accumulation of scalar particles around the black hole, giving rise to observable effects such as alterations in gravitational waveforms and interference with the imaging of the black hole shadow [54, 55, 56, 57, 58, 59, 60]. These observable effects offer valuable insights and avenues for exploring exotic particles beyond the standard model and furthering our comprehension of the physical characteristics of black holes. See more discussions on this phenomenon and its astrophysical implications, please refer to the review [61, 62] and references therein. Thus, the research findings in Refs. [35, 36, 37] not only unveil the superradiant instability of ABG black holes under specific circumstances but also present a potential observational approach for identifying and investigating the properties of regular black holes.